# Relation of Continuous Chirality Measure to Spin and Orbital Polarization, and Chiroptical Properties in Solids

### v1

@@ Abstract
Chirality introduces intriguing topological, electronic, and optical properties to molecules and solids. In this work, we investigate the influence of structural chirality on spin and orbital polarization as well as optical activity through first-principles calculations and continuous chirality measure (CCM). By using chiral selenium (Se) and 2D hybrid perovskites as examples, we demonstrate that chirality continuously modifies spin-orbit splitting and orbital angular momentum (OAM) polarization. We establish a direct relation between chirality transfer across organic-inorganic interfaces and inversion symmetry breaking, which induces Rashba-Dresselhaus spin splitting in hybrid perovskites. Additionally, we examine the effects of chirality on circular dichroism (CD) and the circular photogalvanic effect (CPGE), demonstrating how the continuous tuning of chirality dictates their magnitude and anisotropy. Our findings highlight that applying hydrostatic pressure can effectively tune the CCM, thereby enhancing the chirality-induced effects such as spin selectivity. By integrating CCM with electronic structure calculations, we present a predictive strategy for designing chiral materials with tailored optical and spintronic functionalities.
@@ I Introduction
Structural chirality lacks spatial inversion and mirror symmetries, resulting in distinct electronic and optical responses [1]. This inherent asymmetry plays a crucial role in optical orientation, where chiral materials interact differently with left- and right-circularly polarized light. Such interactions give rise to circular dichroism (CD) [2] and circularly polarized luminescence (CPL) [3, 4]. The ability to control light-matter interactions through chirality enhances spin-dependent optical functionalities in modern optoelectronic applications. Besides optical activity, chirality has important implications in quantum transport; for example large spin polarization can be generated by running charges in nonmagnetic low-spin-orbit chiral materials, known as the chirality-induced spin selectivity (CISS) effect [5, 6]. As such effect implies spontaneous time-reversal symmetry broken in absence of permanent magnetization, the mechanism behind the CISS effect is still debated. Since its discovery in 2012 [5], researchers have proposed different theories, but no clear agreement has been reached [7, 8, 9]. Additionally, chirality plays a key role in orbitronics, an emerging field that utilizes electron orbital angular momentum (OAM) [10, 11, 12, 13] as an information carrier. The broken inversion symmetry in chiral materials enables coupling between OAM and spin angular momentum (SAM) [14, 15, 16, 17], leading to chiral orbital textures and potentially large orbital Hall conductivity (OHC) [13, 18]. By integrating these effects, chiral materials serve as a versatile platform for next-generation optoelectronic, spintronic, and orbitronic applications, enabling advances in quantum information science and low-power electronics [19, 20].
@@
Quantifying chirality is critical to understanding its effects on electronic and optical properties. The Continuous Chirality Measure (CCM) offers a reliable way to do so. Unlike binary classifications of chirality, CCM provides a continuous scale to evaluate the degree of chirality in a given structure. For example, Hausdorff distances and CCM have been applied to crystal structures to distinguish chiral space groups and analyze their influence on band structures and Berry curvatures [21]. In another approach, graph-theoretical chirality (GTC) [22], based on torsion in differential geometry, has been proposed to describe chirality across different length scales, offering direct correlations with optical activity and chiral meta-materials. Beyond structural characterization, CCM has emerged as a crucial parameter in describing the electronic properties of chiral materials, in particular to their spin and orbital textures, as we will show in this work. In systems with the presence of CISS, a larger CCM often correlates with stronger spin-polarization (or spin-splitting in k space), suggesting a direct impact on spin-selective transport. Similarly, in chiral crystals, OAM textures emerge, and as we will show later, their properties can be quantitatively related to CCM. Thus, CCM not only measures chirality geometrically but also predicts emergent quantum properties.
CCM plays a crucial role in linking structural chirality to chiroptical properties such as CD and the Circular Photogalvanic Effect (CPGE). Since CD reflects differential absorption of left- and right-circularly polarized light, higher CCM intuitively correlates with a stronger CD signal. This is because a higher CCM may correspond to more pronounced chiral electronic states or electronic transition properties, thereby enhancing the optical response. Recent studies on chiral perovskites demonstrate that tuning molecular-level chirality through structural engineering enhances CD, reinforcing the connection between CCM and chiroptical properties [23]. Although CD is widely used to characterize chirality in molecules and materials, CPGE has been used to probe the topological charge of Weyl points in chiral materials [24, 25], therefore its precise relationship with CCM needs further experimental and theoretical investigation. On the other hand, CPGE offers a unique probe of chirality-sensitive charge dynamics, as its response is inherently linked to broken crystal symmetry and spin-orbit coupling (SOC) effects such as Rashba and Dresselhaus interactions [26, 27]. Since chirality breaks inversion symmetry, these SOC-induced band splittings can lead to spin-polarized photocurrents under circularly polarized light. This makes CPGE a promising tool for characterizing chiral materials, providing complementary insights beyond traditional chiroptical measurements like CD, particularly in chiral semiconductors and hybrid 2D-Perovskites [27].
@@
The remainder of this paper is organized as follows. In Section II.A, we introduce the definition of the CCM for solid-state crystals and two-dimensional hybrid perovskites, establishing a quantitative framework for measuring the degree of chirality and exploring its correlation with spin and orbital textures in these systems. In Section II.B, we investigate the influence of chirality on optical responses, specifically CD and absorption dissymmetry factor ($g_{CD}$). We discuss how symmetry considerations, including inversion and mirror symmetry, govern these optical effects and further examine the role of strain in modulating CD. In Section II.C, we analyze the relationship between chirality and CPGE, highlighting the impact of SOC and Rashba-Dresselhaus effects on the observed trends. In Section II.D, we explore the transfer of chirality across organic-inorganic interfaces in hybrid perovskites, emphasizing how this process modifies electronic and optical properties, particularly in enhancing CD and CPGE. Finally, in Section III, we summarize our findings and propose future directions for leveraging chirality in the design of next-generation optoelectronic and spintronic materials.
@@ II.1 Continuous Chirality Measure
The CCM has been commonly used to quantify chirality in molecules and was recently adapted to describe periodic crystals [21]. CCM is defined by the mean-squared distance between atomic positions in the chiral structure and their closest equivalent positions in a symmetric reference structure. This approach ensures that the measure effectively captures the degree of chirality. The unnormalized chirality measure is defined as:
@@
where $p_{i}$ are the atomic positions in the target chiral structure, and $p_{i}^{\text{sym}}$ are the nearest equivalent positions in an achiral reference space group such that $\tilde{S}^{2}(G)$ is minimized.
@@
where $N$ is a normalization factor accounting for the difference between the achiral and chiral structure:

### v2

@@ Abstract
Chirality introduces intriguing topological, electronic, and spin-optronic properties to molecules and solids. In this work, we provide a quantitative metric for the degree of chirality in solids, independent of the type of system and the dimensionality, through the continuous chirality measure (CCM). We quantitatively analyze the correlation between CCM and spin and orbital angular momentum (OAM) polarization, as well as circular dichroism (CD) and the circular photogalvanic effect (CPGE). By internal spin-orbit field analysis, we demonstrate a distinct character (proportionality among Rashba, Deresselhaus, and Weyl contributions) and chirality dependence among different chiral solids. Furthermore, unlike CD, we found that absorption dissymmetry factor $g_{CD}$ could remain unchanged as a function of chirality and show anisotropic dependence on CCM. In addition, we show that the relation between CCM and CPGE is rather complex. At low excitation energy close to the bandgap transition, the CCM continuously tunes the total SOC, and therefore, the CPGE response. However, at high excitation energy, CPGE includes more than just band edge transitions, which complicates the relation of chirality and CPGE due to changes in the optical dipole strength and electron-hole group velocity difference. Ultimately, this causes CPGE to be only correlated with chirality at excitation energies close to the band edge. At the end, we discussed strategies of manipulating chiral-optical properties through chirality transfer at interfaces or applying strain. The insights developed in this work will inspire the design of materials for future spintronics and orbitronics, as well as spin-optronics applications.
@@ I Introduction
Structural chirality lacks spatial inversion and mirror symmetries, resulting in distinct electronic and optical responses [siegel1998homochiral]. This inherent asymmetry plays a crucial role in optical orientation, where chiral materials interact differently with left- and right-circularly polarized light. Such interactions give rise to circular dichroism (CD) [kelly2005study] and circularly polarized luminescence (CPL) [mun2020electromagnetic, sang2020circularly]. The ability to control light-matter interactions through chirality enhances spin-dependent optical functionalities in modern optoelectronic applications. Besides optical activity, chirality has important implications in spin transport; for example large spin polarization can be generated by running charges in nonmagnetic low-spin-orbit chiral materials, known as the chirality-induced spin selectivity (CISS) effect [naaman2012chiral, naaman2019chiral]. Such effect implies spontaneous time-reversal symmetry broken in absence of permanent magnetization. Since its discovery in 2012 [naaman2012chiral], researchers have proposed different theories to explain its mechanism, but no clear agreement has been reached [calavalle2022gate, roy2022long, tenzin2023collinear]. Additionally, chirality plays a key role in orbitronics, an emerging field that utilizes electron orbital angular momentum (OAM) [liu2021chirality, kim2023optoelectronic, yang2023monopole, brinkman2024chirality] as an information carrier. The broken inversion symmetry in chiral materials enables coupling between OAM and spin angular momentum (SAM) [hirayama2015weyl, sakano2020radial, gatti2020radial, yang2025chirality], leading to chiral orbital textures and potentially large orbital Hall conductivity (OHC) [brinkman2024chirality, go2018intrinsic]. By integrating these effects, chiral materials serve as a versatile platform for next-generation optoelectronic, spintronic, and orbitronic applications, enabling advances in quantum information science and low-power electronics [yang2021chiral, forment2022chiral].
@@
Quantifying chirality is critical to understanding and controlling its effects on electronic and optical properties. The Continuous Chirality Measure (CCM) offers a reliable way to do so. Unlike binary classifications of chirality, CCM provides a continuous scale to evaluate the degree of chirality in a given structure. For example, Hausdorff distances and CCM have been applied to crystal structures to distinguish chiral space groups and analyze their influence on band structures and Berry curvatures [fecher2022chirality]. In another approach, graph-theoretical chirality (GTC) [cha2024graph], based on torsion in differential geometry, has been proposed to describe chirality across different length scales, offering direct correlations with optical activity and chiral meta-materials. Beyond structural characterization, CCM has emerged as a crucial parameter in describing the electronic properties of chiral materials, in particular to their spin and orbital textures, as we will show in this work. In systems with the presence of CISS, a larger CCM often correlates with stronger spin-polarization (or spin-splitting in k space), suggesting a direct impact on spin-selective transport. Similarly, in chiral crystals, OAM textures emerge and their properties can be quantitatively related to CCM, as shown in this work. Thus, CCM not only measures chirality geometrically but also predicts emergent quantum properties [Chang2018-jj].
CCM plays a crucial role in linking structural chirality to chiroptical properties such as CD and the Circular Photogalvanic Effect (CPGE). Since CD reflects differential absorption of left- and right-circularly polarized light, higher CCM intuitively correlates with a stronger CD signal. This is because a higher CCM may correspond to more pronounced chiral electronic states or electronic transition properties, thereby enhancing the optical activity. Experimentally, the absorption dissymmetry factor ($g_{CD}$) is often used to characterize the degree of chirality, which normalizes the CD by the optical absorption. However, the correlation between $g_{CD}$ and the degree of chirality remains to be validated, and is examined in this work. Recent studies demonstrate that tuning chiral molecule-surface interactions can realize chirality transfer to nonchiral surfaces and enhances CD, reinforcing the connection between CCM and chiroptical properties [Haque2024-yw, song2024enhancing, apergi2023calculating].
Besides CD, CPGE is another common optical characterization technique for chiral materials. With circularly-polarized light excitation, DC photocurrent is generated in chiral solids, and governed by characteristic symmetry relations [le2021topology]. It also has been used to probe the topological charge of Weyl points in chiral materials [de2017quantized, le2020ab]. Furthermore, CPGE offers a unique probe of chirality-sensitive charge dynamics, as its response is inherently linked to broken crystal symmetry and spin-orbit coupling (SOC) effects [yu2015temperature, zhu2024electrically]. Since chirality breaks inversion symmetry, these SOC-induced band splittings can lead to spin-polarized photocurrents under circularly-polarized light. This makes CPGE a promising tool for characterizing chiral materials, providing complementary insights beyond traditional chiroptical measurements like CD, particularly in chiral semiconductors and 2D hybrid organic-inorganic perovskites (2D HOIP) [zhu2024electrically]. Despite its importance, the relation between CPGE and the degree of chirality remains elusive and requires in-depth investigations.
Previous works have investigated the correlation between structural properties and SOC spin-splitting as well as CD in 2D-HOIPs. These studies have focused on structural parameters unique to halide perovskites [jana2021structural, chakraborty2024design, fortino2024role, pols2024temperature, apergi2023calculating]. For example, one measure is the $\Delta\beta_{in}$, which is the difference in M-X-M bond angles projected to the plane of the 2D inorganic sublattice [jana2021structural]. They find a correlation between $\Delta\beta_{in}$ and the Rashba parameters, but a later study found $\Delta\beta_{in}$ is not suitable for Aba2 and Pbcn space-groups, where $\Delta\beta_{in}=0$ despite large spin splittings [chakraborty2024design]. In another study, CD was computed from the tight-binding theory and did not correlate with $\Delta\beta_{in}$ either [apergi2023calculating]. Instead they correlated CD with the relative distances of neighboring axial halide atoms with the center atom of the octahedral cage. Other studies have used ab initio molecular dynamics (AIMD) to investigate temperature effects on chiral perovskites [pols2024temperature, fortino2024role]. They found a reduction in chirality as temperature increases [pols2024temperature]. Additionally, tin-based perovskites have a larger distortion due to weaker bond affinity  [fortino2024role]. While past works focus on specific structural properties of 2D-HOIPs, here we focus on using a universal metric for chiral molecules and solids, i.e. continuous-chirality measure, to investigate its relation to spin and chiral-optical properties across different classes of materials. To this end we choose two disparate prototypical chiral systems: one 2D HOIP (2D-NPB) and one inorganic single-element solid, Se. Instead of including a large number of systems, we aim for an in-depth analysis of these two materials, which have distinct symmetry, chemical bonding, potentially representative of a larger class of materials. Additionally they have comprehensive experimental studies for theoretical validation. In contrast to previous studies, we quantify the relation of the degree of chirality with a more comprehensive set of observables, e.g. including CPGE and OAM for the first time.
In addition to inversion symmetry breaking, perovskites are known to have dynamic disorder that induces an effect such as dynamical Rashba  [li2024spin, abramovitch_thermal_2021]. Specifically, in the centrosymmetric perovskite MAPbBr3, the upper limit of dynamical Rashba splitting was found to be 1.35 eV/Å [li2024spin]. This indicates the significance of this effect in perovskites. While this is indeed an area of research that warrants further attention, it is outside the scope of this work. Instead, we focus on the effects of chirality alone. Additionally, we find excellent agreement with experimental CD for 2D-NPB (except for the missing excitonic peak below the bandgap transitions); see Figure 8d. Therefore, we believe that the contribution from chirality-induced symmetry breaking significantly outweighs the effects from dynamical disorder.
@@
In this work we present a quantitative assessment of relation between the degree of chirality (CCM) and a comprehensive set of spin-optotronic properties (CD, $g_{CD}$, CPGE, OAM, and spin) across chiral materials. The remainder of this paper is organized as follows. In Section 2.1, we introduce the definition of CCM for solid-state crystals and two-dimensional hybrid perovskites, establishing a quantitative framework for measuring the degree of chirality and exploring its correlation with spin and orbital textures in these systems. In Section 2.2, we investigate the influence of chirality on optical responses, specifically CD and absorption dissymmetry factor ($g_{CD}$). We discuss the anisotropic dependence of $g_{CD}$ on chirality and further examine the role of strain in modulating CD. In Section 2.3, we analyze the relationship between chirality and CPGE, highlighting the complex factors determining CPGE beyond SOC spin-splitting. In Section 2.4, we explore the transfer of chirality across organic-inorganic interfaces in hybrid perovskites, emphasizing how this process modifies both electronic and optical properties, particularly in enhancing CD and CPGE. Finally, in Section 3, we summarize our findings and propose future directions for leveraging chirality in the design of next-generation optoelectronic, spintronic and orbitronic materials.
@@ II.1 Continuous Chirality Measure
The CCM has been commonly used to quantify the chirality in molecules and was recently adapted to describe periodic crystals [fecher2022chirality]. CCM is defined by the mean squared distance between atomic positions in the chiral structure and their closest equivalent positions in a symmetric reference structure. This approach ensures that the measure effectively captures the degree of chirality. The unnormalized chirality measure is defined as:
@@
where $p_{i}$ are the atomic positions in the target chiral structure and $p_{i}^{\text{sym}}$ are the closest equivalent positions in an achiral reference space group ($G$) such that $\tilde{S}^{2}(G)$ is minimized.
@@
where $N$ is a normalization factor that accounts for the difference between the achiral and chiral structures: